OpenStudy (anonymous):

If the product xy = 28 when y = 196, what is the value of x? Express your answer as a common fraction.

OpenStudy (idealist10):

xy=28 y=196 x*196=28 x=28/196

OpenStudy (idealist10):

if your teacher wants the fraction to be simplified, then x=1/7.

OpenStudy (anonymous):

thanks so much!

OpenStudy (idealist10):

You're welcome.